# Survey on Hand Gesture Recognition from Visual Input

### v1

@@ IV-E Summary of Methodology Overview
To provide a clear understanding of the groupings, Table III summarizes the classification of the articles based on the criteria described in this section. Specifically, these criteria include:
Input Input type Capture Task Method Article RGB-D monocular box/filter classification hybrid [119] skeleton estimation hybrid [173, 98] RGB-D+RGB_image monocular box/filter classification NN [172] classification hybrid [26, 137] estimation hybrid [49] skeleton estimation hybrid [15, 94] RGB-D+video monocular box/filter classification NN [86] classification hybrid [78, 22, 40, 195] estimation non-NN [158] estimation hybrid [147, 159] skeleton classification NN [149, 123] classification hybrid [84, 11, 162, 59, 90] classification non-NN [31] classification+estimation NN [41] RGB_image monocular box/filter classification NN [188, 4, 24, 148, 21] classification hybrid [12, 106, 144, 196, 97, 82] classification non-NN [10, 38] estimation NN [71, 143] estimation hybrid [130, 145, 76, 75, 95, 60] skeleton classification NN [153] classification hybrid [171] classification+estimation hybrid [9] estimation NN [56, 126, 168, 191] estimation hybrid [129, 107, 151] multiview skeleton estimation hybrid [198] Video monocular box/filter classification NN [114, 87, 121, 69, 160] classification hybrid [139, 73, 74, 77, 101, 199, 58, 115, 128, 81, 79, 67, 182, 28, 104, 180, 89, 6, 36, 154, 42, 39, 29, 136] estimation NN [176] estimation hybrid [72, 37, 184] skeleton classification NN [34, 33, 117, 45] classification hybrid [1, 120, 65, 164, 51, 200, 105, 61] classification+estimation hybrid [186, 111] classification+estimation non-NN [66] estimation NN [112, 16] estimation hybrid [166] multiview box/filter classification non-NN [133] skeleton classification hybrid [183] estimation NN [23] estimation hybrid [64]
TABLE III: Summary of HGR Methods based on classification criteria of Section IV.
@@
Combinations of values within a specific group/column are indicated using the ”+” character. For instance, an entry of RGB-D+video in the ”Input Type” column signifies that the referenced articles utilize a combination of RGB-D and video data for hand gesture recognition. This notation allows for a concise representation of articles employing multiple values of a group.
@@ IV-F Association between input, input type, capture and hand recognition task
The results are visualized in Figures 11 to 14, where the conditional probabilities $P(Y|X)$ are shown for each combination of X and Y. Each subplot corresponds to a specific group, with blue bars representing the classification goal and red bars representing the estimation goal. Like so, the plots provide valuable insights into the relationship between different groups and the probability of achieving classification or estimation goals.
@@ V-A Support Vector Machines (SVMs)
This function outputs $f(\mathbf{x}_{new})=y_{j}$, where $j=1,2,...,C$, indicating which class the new gesture belongs to.
@@ V-B Hidden Markov Models (HMMs)
  • Transition Probabilities: To model the transitions between these hidden states, transition probabilities are defined. For example, let $P(s_{t}=j|s_{t-1}=i)=a_{ij}$ denote the probability of transitioning from state $i$ at time $t-1$ to state $j$ at time $t$. This probabilistic framework allows us to account for the inherent variability in how gestures are performed.
@@ V-C Convolutional Neural Networks (CNNs)
In the gesture recognition task using CNNs, each image is first represented as a 3D tensor. Let =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{N}\}$ be the set of input images representing different gestures, where $N$ is the total number of images. Each image $I_{i}$ can be represented as a 3D tensor $I_{i}\in\mathbb{R}^{H\times W\times C}$, where $H,W$ are the image height and width, respectively, and $C$ is the number of color channels (e.g., RGB).
@@ V-E Graph Convolutional Networks (GCNs)
The relationships between nodes are then described using an adjacency matrix $A$, where $A_{ij}=1$ if there is an edge between nodes $v_{i}$ and $v_{j}$ and 00 otherwise.
@@
where $H^{(l)}$ is the matrix of node features at layer $l$, $W^{(l)}$ is the weight matrix for layer $l$, $A$ is the adjacency matrix, $\sigma(\cdot)$ is an activation function (e.g., ReLU).
@@ V-F Transformer Models
The introduction of transformer-based models has significantly advanced HGR, offering superior performance in tasks requiring the capture of long-range dependencies and intricate spatial-temporal relationships [191, 51, 186].
@@
Innovations like hand-level tokenization reduce computational complexity while maintaining high accuracy. Multi-task transformer networks, designed for simultaneous gesture recognition and pose estimation, have further optimized performance through collaborative learning. These models excel in scenarios involving hand-object interactions or continuous sign language translation, achieving competitive results on benchmark datasets. Their ability to encode contextual features with precision highlights their transformative impact on HGR.
VI Datasets
@@
This section categorizes datasets based on various criteria, including input type (e.g., RGB, RGB-D, video), number of samples, resolution, number of classes (for classification datasets), and gesture type (dynamic or static). Additionally, the categorization incorporates the primary hand gesture recognition tasks discussed in Section II-D. Since the same dataset can be used for multiple tasks and can be employed in multiple research works, we determine the primary task associated with each dataset, by considering the task that it is applied to in the majority of articles that use the dataset. For instance, if most articles that use a dataset, use it for Hand Gesture Classification and only a few of them use it for Sign Language Recognition, the dataset is categorized under Hand Gesture Classification.
@@
The datasets used for classification tasks such as Hand Gesture Classification and Sign Language Recognition are presented in Table IV. In a similar manner, Table V displays the datasets used for estimation tasks, including Hand Gesture Estimation and Hand/Body Reconstruction.
@@
TABLE IV: Overview of datasets for hand gesture classification and sign language recognition.
@@
Nigerian Sign Language [70] 2015 25900 images Not specified 37 BdSL [50] 2020 26713 images 300x500 pixels 36 PSL [54] 2021 1480 images 480x640 pixels 37 Japanese Sign Language [106] 2021 7380 images 400x400 pixels 41 Arabic Alphabet Sign Language [3] 2023 7534 images Not specified 31 RGB+RGB-D American Sign Language [134] 2011 48000 samples 128x128 pixels 24 Dynamic RGB LSA64 [142] 2016 3200 videos 1920x1080 pixels 64 Korean Sign Language [185] 2019 1540 videos Not specified 77 British Sign Language [1] 2020 16497 frames Not specified 18 BosphorusSign22k [124] 2020 22542 videos Not specified 744 Word-Level American Sign Language [88] 2020 34404 videos Not specified 2000 Indian Sign Language [155] 2020 4287 videos Not specified 263 MINDS-Libras (BRASL) [140] 2021 1200 videos 1080x1920 pixels 20 Chinese sign language [182] 2024 25000 videos Not specified 178 RGB-D Greek Sign Language [1] 2012 4910 videos Not specified 20 RGB-D+RGB+Skeleton AUTSL [152] 2020 38336 videos 512x512 pixels 226 video PHOENIX-2014 [36] 2014 6841 videos Not specified 1295 Microsoft American Sign Language [65] 2018 25000 videos Not specified 1000 RWTH-PHOENIX14T [17] 2018 0.95 million frames Not specified 3953 ALLVD [27] 2019 7798 videos Not specified 2745
@@
TABLE V: Overview of datasets for hand gesture estimation and hand/body reconstruction.
@@
Topic Type of Gesture Input Dataset Year # Samples Resolution Hand Gesture Estimation Static RGB MPII+NZSL [56] 2018 2800 images Not specified RGB OneHand10K [175] 2019 11703 images Not specified RGB FreiHAND [198] 2019 37k frames 224x224 pixels RGB+RGB-D Stereo Hand [189] 2016 18000 RGB+18000 RGB-D Not specified RGB+RGB-D ObMan [9] 2018 141550 images 256x256 pixels RGB-D ICVL [161] 2014 180K images Not specified Dynamic RGB Rendered Hand Pose [197] 2017 43986 images 320x320 pixels BiliHands [176] 2024 11204 images Not specified GANerated [112] 2018 331k frames 256x256 pixels InterHand2.6M [110] 2020 2590k frames 512x334 pixels RGB+RGB-D HIC [167] 2016 29 sequences 640x480 pixels B2RGB [125] 2017 1888 frames Not specified SHOWMe [159] 2023 87540 frames Not specified RGB-D Dexter+Object [156] 2016 3000 frames 640x480 pixels EgoDexter [113] 2017 3k frames 640x480 pixels Hands2017 [187] 2017 1.2 million frames 640x480 pixels SynthHands [113] 2017 220k frames 640x480 pixels HO-3D [47] 2020 78k frames 640x480 pixels DexYCB [20] 2021 582k frames 640x480 pixels RGB-D+video FORTH [122] 2011 7148 frames Not specified ASTAR [181] 2013 435 frames Not specified Dexter [157] 2013 3157 frames Not specified UCI-EGO [141] 2014 3640 frames Not specified NYU [163] 2014 8252 frames Not specified MSRA [135] 2014 2400 frames Not specified ICL [161] 2014 1599 Not specified Hand/Body Reconstruction Static RGB LSP [62] 2010 2000 images Not specified COCO [92] 2014 328000 images Not specified Expressive hands and faces dataset [129] 2019 100 images Not specified LSP-Extended [62] 2019 10000 images Not specified Dynamic RGB MPI-INF-3DHP [103] 2017 1.3 million frames Not specified Kinetics-400 [19] 2017 400 videos Not specified InstaVariety [68] 2019 2.1 million frames Not specified RGB+3D Human3.6M [55] 2013 3.6 million images 50Hz video RGB+Skeleton 3DPW [169] 2018 60 videos Not specified RGB+RGB-D+video CMU [63] 2015 297000 frames Not specified video PennAction [192] 2013 2326 videos Not specified PoseTrack [7] 2018 1337 videos Not specified
@@ VI-A Dataset statistics
Figures 16 and 17 illustrate the distribution of datasets analyzed in this study. Specifically, Fig. 16 highlights that most datasets were created in 2018. The darker hues in the plot indicate that the most frequently used datasets were created in 2020, followed by those from 2016 and 2018. Overall, datasets created between 2016 and 2020 are the most utilized, likely because they strike a balance between recency and maturity. Older datasets (e.g., 2008–2014) may be less relevant due to outdated methods (e.g., lower resolution or fewer samples), while more recent datasets (2021 and later) may not yet have been widely adopted or refined.

### v3

@@ IV-E Summary of Methodology Overview
To provide a clear understanding of the groupings, Table IV summarizes the classification of the articles based on the criteria described in this section. Specifically, these criteria include:
TABLE IV: Summary of HGR Methods based on classification criteria of Section IV.
Input Input type Capture Task Method Article RGB-D monocular box/filter classification NN [143, 48, 230] skeleton estimation hybrid [211, 121] RGB-D+RGB_image monocular box/filter classification NN [210] classification hybrid [34, 165] estimation hybrid [66] skeleton estimation hybrid [20, 115] RGB-D+video monocular box/filter classification NN [106] classification hybrid [98, 29, 51, 237] estimation non-NN [192] estimation hybrid [179, 193] skeleton classification NN [182, 148, 113] classification hybrid [104, 15, 197, 79, 109] classification non-NN [39] classification+estimation NN [52] RGB_image monocular box/filter classification NN [229, 7, 31, 181, 28, 3] classification hybrid [16, 130, 174, 238, 120, 102, 14] classification non-NN [13, 46] estimation NN [91, 173] estimation hybrid [156, 175, 96, 95, 117, 80] skeleton classification NN [186] classification hybrid [209, 6] classification+estimation hybrid [12] estimation NN [75, 152, 205, 233, 111, 116, 188] estimation hybrid [155, 131, 184] multiview skeleton estimation hybrid [240] Video monocular box/filter classification NN [138, 107, 145, 89, 194, 49, 217] classification hybrid [167, 93, 94, 97, 124, 241, 78, 139, 154, 101, 99, 87, 221, 36, 128, 219, 108, 9, 44, 187, 55, 47, 37, 164] estimation NN [214] estimation hybrid [92, 45, 223] skeleton classification NN [42, 41, 141, 60, 58] classification hybrid [1, 144, 85, 199, 69, 242, 129, 81] classification+estimation hybrid [226, 135] classification+estimation non-NN [86] estimation NN [136, 21] estimation hybrid [202] multiview box/filter classification non-NN [160] skeleton classification hybrid [222] estimation NN [30, 59] estimation hybrid [84]
@@
Combinations of values within a specific group/column are indicated using the "+" character. For instance, an entry of RGB-D+video in the "Input Type" column signifies that the referenced articles utilize a combination of RGB-D and video data for hand gesture recognition. This notation allows for a concise representation of articles employing multiple values of a group.
@@ IV-F Association between input, input type, capture and hand recognition task
The results are visualized in figures 13 to 16, where the conditional probabilities $P(Y|X)$ are shown for each combination of X and Y. Each subplot corresponds to a specific group, with blue bars representing the classification goal and red bars representing the estimation goal. Like so, the plots provide valuable insights into the relationship between different groups and the probability of achieving classification or estimation goals.
@@
IV-G Key Takeaways of this section
This section provided a comprehensive classification of Hand Gesture Recognition methods based on several key factors. The primary ways HGR methods are categorized and their prevalence are:
  • Task Objective: HGR methods are primarily categorized by their goal: Hand Gesture Classification (identifying semantic meaning) or Hand Gesture Estimation (representing hand topology/structure).
  • Input Data Type: Methods utilize different visual inputs including RGB, RGB-D, and Video, with video-based methods being most prevalent.
  • Camera Setup: Monocular camera setups dominate due to their simplicity and cost-effectiveness over multi-view setups.
  • Hand Gesture Capture: Techniques are broadly classified into Skeleton-Based captures (representing the hand as key joints and connections) and Box/Filter-Based captures (detecting the hand regions).
  • Recognition Techniques: Predominantly Hybrid Approaches, combining strengths of multiple techniques (e.g., CNN+LSTM), followed by standalone Neural Network approaches and fewer Non-Neural Network methods.
  • Observed Associations: Analysis showed tendencies such as box/filter methods being more frequently used for classification tasks, while multi-view setups are more common for estimation tasks. Overall, classification is the more common goal across most analyzed categories.
@@ V-A Support Vector Machines (SVMs)
This function outputs $f(\mathbf{x}_{new})=y_{j}$, where j=1,2,,Cj=1,2,...,C, indicating which class the new gesture belongs to.
@@ V-B Hidden Markov Models (HMMs)
  • Transition Probabilities: To model the transitions between these hidden states, transition probabilities are defined. For example, let $P(s_{t}=j|s_{t-1}=i)=a_{ij}$ denote the probability of transitioning from state $i$ at time t1t-1 to state $j$ at time $t$. This probabilistic framework allows us to account for the inherent variability in how gestures are performed.
@@ V-C Convolutional Neural Networks (CNNs)
In the gesture recognition task using CNNs, each image is first represented as a 3D tensor. Let $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{N}\}$ be the set of input images representing different gestures, where $N$ is the total number of images. Each image $I_{i}$ can be represented as a 3D tensor $I_{i}\in\mathbb{R}^{H\times W\times C}$, where $H,W$ are the image height and width, respectively, and $C$ is the number of color channels (e.g., RGB).
@@ V-E Graph Convolutional Networks (GCNs)
The relationships between nodes are then described using an adjacency matrix $A$, where $A_{ij}=1$ if there is an edge between nodes $v_{i}$ and $v_{j}$ and 0 otherwise.
@@
where $H^{(l)}$ is the matrix of node features at layer $l$, $W^{(l)}$ is the weight matrix for layer $l$, AA is the adjacency matrix, $\sigma(\cdot)$ is an activation function (e.g., ReLU).
@@ V-F Transformer Models
The introduction of transformer-based models has significantly advanced HGR, offering superior performance in tasks requiring the capture of long-range dependencies and intricate spatial-temporal relationships [233, 69, 226]. Transformer-based architectures have emerged as powerful tools since they can address both static and dynamic gesture tasks. For instance, GestFormer [54] introduced multiscale wavelet pooling to better model spatiotemporal variations in dynamic gestures, addressing challenges like hand orientation and trajectory changes. Meanwhile, HGR-ViT [195] adapted Vision Transformers (ViTs) for static gesture recognition, achieving robust performance on ASL and NUS datasets by encoding hand orientation and positional features. Recent advancements also emphasize efficiency: ConvMixFormer [53] reduced parameter counts while maintaining generalization for dynamic gestures.
@@
The primary advantage of transformers in HGR lies in their exceptional ability to model long-range dependencies and to process sequences in parallel, leading to potentially faster training and a more holistic understanding of gestures compared to sequential models like LSTMs. They also offer flexibility for transfer learning and multimodal fusion. However, transformers are not without drawbacks. Their standard self-attention mechanism exhibits computational complexity quadratic to the input sequence length ($O(N^{2})$), posing challenges for very long video sequences or high-resolution inputs. This can impact inference speed, particularly on resource-constrained devices. Furthermore, transformers typically require substantial amounts of training data, due to their weaker inductive biases compared to CNNs, and can be harder to interpret, despite the insights offered by attention maps. The necessity for positional encodings to manage their permutation-equivariant nature also adds a layer of design consideration.
When compared to established CNN+LSTM stacks, transformers offer a different paradigm for temporal modeling. While LSTMs inherently capture temporal order and excel at short-to-medium range dependencies, they can struggle with very long sequences. Transformers, in contrast, are theoretically better suited for such long-range interactions due to their global attention, but their $O(N^{2})$ complexity can be more demanding than the LSTMs’ $O(N)$ per-step one for extremely long sequences. CNNs remain highly effective for initial spatial feature extraction due to strong inductive biases, and many HGR transformers leverage CNN backbones. Thus, transformers might be less data-efficient than CNN+LSTM approaches on smaller datasets.
@@
Innovations like hand-level tokenization reduce computational complexity while maintaining high accuracy. Multi-task transformer networks, designed for simultaneous gesture recognition and pose estimation, have further optimized performance through collaborative learning. These models excel in scenarios involving hand-object interactions or continuous sign language translation, achieving competitive results on benchmark datasets. Their ability to encode contextual features with precision highlights their transformative impact on HGR. These developments align with findings from Hashi et al. [63], whose systematic review highlights transformers’ growing dominance in vision-based HGR since 2018, citing improvements in accuracy and computational efficiency. Together, these works illustrate how transformers address diverse HGR challenges, from temporal modeling to cross-dataset generalization.
V-G Key Takeaways of this section
This section explored the prominent algorithms utilized in HGR, highlighting their unique strengths in modeling spatial and temporal aspects of hand gestures. The main algorithms discussed include:
  • Support Vector Machines (SVMs): Remain effective for gesture classification, often paired with handcrafted features like HOG or LEMMP, and can be extended with kernel methods for dynamic gestures.
  • Hidden Markov Models (HMMs): Widely employed for modeling the sequential nature of dynamic gestures, particularly in sign language recognition, by representing gestures as sequences of hidden states.
  • Convolutional Neural Networks (CNNs): A cornerstone in HGR for their ability to automatically learn and extract hierarchical spatial features directly from visual data (images/video frames).
  • Long Short-Term Memory (LSTM) Networks: A type of Recurrent Neural Network (RNN) excelling at capturing temporal dependencies in sequential gesture data, frequently combined with CNNs for spatio-temporal modeling.
  • Graph Convolutional Networks (GCNs): Suited for HGR tasks where data can be represented as a graph (e.g., hand skeletons), allowing for the modeling of non-Euclidean structural relationships between joints.
  • Transformer Models: Have shown significant advancements in HGR, effectively capturing long-range dependencies and complex spatio-temporal patterns in both static and dynamic gestures, leading to state-of-the-art performance.
VI Datasets and metrics
@@
This section presents the main datasets used as benchmarks in HGR tasks and gives an overview of the main evaluation metrics. It categorizes datasets based on various criteria, including input type (e.g., RGB, RGB-D, video), number of samples, resolution, number of classes (for classification datasets), and gesture type (dynamic or static). Additionally, the categorization incorporates the primary hand gesture recognition tasks discussed in Section II-D. Since the same dataset can be used for multiple tasks and can be employed in multiple research works, we determine the primary task associated with each dataset, by considering the task that it is applied to in the majority of articles that use the dataset. For instance, if most articles that use a dataset, use it for Hand Gesture Classification and only a few of them use it for Sign Language Recognition, the dataset is categorized under Hand Gesture Classification.
@@
The datasets used for classification tasks such as Hand Gesture Classification and Sign Language Recognition are presented in Table V. In a similar manner, Table VI displays the datasets used for estimation tasks, including Hand Gesture Estimation and Hand/Body Reconstruction.
@@
TABLE V: Overview of datasets for hand gesture classification and sign language recognition.
@@
Nigerian Sign Language [90] 2015 25900 images Not specified 37 BdSL [67] 2020 26713 images 300x500 pixels 36 PSL [73] 2021 1480 images 480x640 pixels 37 Japanese Sign Language [130] 2021 7380 images 400x400 pixels 41 Arabic Alphabet Sign Language [4] 2023 7534 images Not specified 31 RGB+RGB-D American Sign Language [162] 2011 48000 samples 128x128 pixels 24 Dynamic RGB LSA64 [172] 2016 3200 videos 1920x1080 pixels 64 Korean Sign Language [224] 2019 1540 videos Not specified 77 British Sign Language [1] 2020 16497 frames Not specified 18 BosphorusSign22k [150] 2020 22542 videos Not specified 744 WLASL [107] 2020 34404 videos Not specified 2000 Indian Sign Language [189] 2020 4287 videos Not specified 263 MINDS-Libras (BRASL) [168] 2021 1200 videos 1080x1920 pixels 20 Chinese sign language [221] 2024 25000 videos Not specified 178 RGB-D Greek Sign Language [1] 2012 4910 videos Not specified 20 RGB-D+RGB+Skeleton AUTSL [185] 2020 38336 videos 512x512 pixels 226 video PHOENIX-2014 [44] 2014 6841 videos Not specified 1295 Microsoft American Sign Language [85] 2018 25000 videos Not specified 1000 RWTH-PHOENIX14T [22] 2018 0.95 million frames Not specified 3953 ALLVD [35] 2019 7798 videos Not specified 2745
@@
TABLE VI: Overview of datasets for hand gesture estimation and hand/body reconstruction.
@@
Topic Type of Gesture Input Dataset Year # Samples Resolution Hand Gesture Estimation Static RGB MPII+NZSL [75] 2018 2800 images Not specified RGB OneHand10K [213] 2019 11703 images Not specified RGB FreiHAND [240] 2019 37k frames 224x224 pixels RGB+RGB-D Stereo Hand [231] 2016 18000 RGB+18000 RGB-D Not specified RGB+RGB-D ObMan [12] 2018 141550 images 256x256 pixels RGB-D ICVL [196] 2014 180K images Not specified Dynamic RGB RHP [239] 2017 43986 images 320x320 pixels BiliHands [214] 2024 11204 images Not specified GANerated [136] 2018 331k frames 256x256 pixels InterHand2.6M [134] 2020 2590k frames 512x334 pixels RGB+RGB-D HIC [203] 2016 29 sequences 640x480 pixels B2RGB [151] 2017 1888 frames Not specified SHOWMe [193] 2023 87540 frames Not specified RGB-D Dexter+Object [190] 2016 3000 frames 640x480 pixels EgoDexter [137] 2017 3k frames 640x480 pixels Hands2017 [227] 2017 1.2 million frames 640x480 pixels SynthHands [137] 2017 220k frames 640x480 pixels HO-3D [62] 2020 78k frames 640x480 pixels DexYCB [27] 2021 582k frames 640x480 pixels RGB-D+video FORTH [146] 2011 7148 frames Not specified ASTAR [220] 2013 435 frames Not specified Dexter [191] 2013 3157 frames Not specified UCI-EGO [171] 2014 3640 frames Not specified NYU [198] 2014 8252 frames Not specified MSRA [163] 2014 2400 frames Not specified ICL [196] 2014 1599 Not specified Hand/Body Reconstruction Static RGB LSP [82] 2010 2000 images Not specified COCO [112] 2014 328000 images Not specified Expressive hands and faces [155] 2019 100 images Not specified LSP-Extended [82] 2019 10000 images Not specified Dynamic RGB MPI-INF-3DHP [127] 2017 1.3 million frames Not specified Kinetics-400 [25] 2017 400 videos Not specified InstaVariety [88] 2019 2.1 million frames Not specified RGB+3D Human3.6M [74] 2013 3.6 million images 50Hz video RGB+Skeleton 3DPW [207] 2018 60 videos Not specified RGB+RGB-D+video CMU [83] 2015 297000 frames Not specified video PennAction [234] 2013 2326 videos Not specified PoseTrack [10] 2018 1337 videos Not specified
@@ VI-A Dataset statistics
Figures 18 and 19 illustrate the distribution of datasets analyzed in this study. Specifically, Fig. 18 highlights that most datasets were created in 2018. The darker hues in the plot indicate that the most frequently used datasets were created in 2020, followed by those from 2016 and 2018. Overall, datasets created between 2016 and 2020 are the most utilized, likely because they strike a balance between recency and maturity. Older datasets (e.g., 2008–2014) may be less relevant due to outdated methods (e.g., lower resolution or fewer samples), while more recent datasets (2021 and later) may not yet have been widely adopted or refined.